OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y – 3z = –8 2x + 2y + z = 12 3x + y – z = –2

OpenStudy (anonymous):

multiply second by 3 and add to give 7x+7x = 28 add second and third 5x + 3y = 10

OpenStudy (anonymous):

These are the answer choices. so it would be the second choice right? 7x + 7y = 28 3x + y = –2 7x + 7y = 28 5x + 3y = 10 x + y = –8 3x + y = –2 x + y = –8 5x + 3y = 14